# Subgroup separability of twisted right-angled Artin groups

### v1

@@ Abstract
We characterise twisted right-angled Artin groups that are subgroup separable using only their defining mixed graphs; we deduce that the group is subgroup separable if and only if the underlying simplicial graph of the mixed graph does not contain paths or squares in 4 vertices. Our work generalises the results of Metaftsis - Raptis on classical right-angled Artin groups.
@@ 1 Introduction
In geometric group theory, the class of right-angled Artin groups (RAAGs shortly) plays a major role, see (Charney, 2007) for a survey. RAAGs are also known as graph groups.
@@
This article deals with the same problem, with focus in another class of groups that generalises the class of RAAGs. This is the class of twisted right-angled Artin groups (T-RAAGs shortly). T-RAAGs appear in (Clancy and Ellis, 2010), where the notion of a twisted Artin group was introduced. We refer to T-RAAGs also with the name twisted graph groups. More details regarding their normal forms and other properties can be found in the author’s PhD thesis (Foniqi, 2022, Chapter 3), and also in (Foniqi, 2024).
@@
Our main work, outlined in Section 4, characterises T-RAAGs that are subgroup separable. This characterisation shows that a T-RAAG is subgroup separable, if and only if its defining mixed graph does not contain paths or squares in $4$ vertices. Here, in contrast with RAAGs, some of the edges in our graphs can be directed. In the case when the graph has no directed edges, we recover the result of Metaftsis and Raptis (2008) for RAAGs.
@@ 2.2 Twisted right-angled Artin groups
To define tRAAGs we use mixed graphs, which are similar to simplicial graphs but directed edges are allowed. We define mixed graphs using simplicial graphs as an underlying structure.
@@
Word problem is one of the fundamental algorithmic questions in algebra; for a group $G$, generated by a set $A$, the word problem asks whether $w\in(A\cup A^{-1})^{\ast}$ is equal to $1_{G}$ in $G$. This problem is solved for T-RAAGs in (Foniqi, 2022) and (Foniqi, 2024).
A more general problem is the subgroup membership problem (know also as the generalised word problem) for a group $G$, which asks for the existence of an algorithm that takes as input an element $g\in G$ and a finitely generated subgroup $H\leqslant G$ and outputs a YES – NO answer whether $g$ belongs to $H$ or not. This is still open for RAAGs, hence for T-RAAGs as well. We will see shortly, that there is a useful connection between being subgroup separable and having decidable subgroup membership problem.
@@ 2.4 Reidemeister-Schreier procedure
then $Y=\{tx(\overline{tx})^{-1}\mid t\in T,\;x\in X,\;tx\not\in T\}$ gives a set of generators for $H$.
@@ 3 Subgroup separability for RAAGs
Summarising the results here, we obtain the other direction of the goal as well, as desired.
@@
In [Mikhailova, 1966] it is shown that $A(C_{4})=F_{2}\times F_{2}$ contains a subgroup where membership is undecidable.
@@ 4 Going from RAAGs to T-RAAGs
We will answer one part of this question by applying the normal forms in T-RAAGs coming from (Foniqi, 2022) and Foniqi (2024), which yield the following two results:
@@
Combining this with the results in Section 3, we obtain:
@@ 4.1 Subgroup Separability for T-RAAGs
By induction, the group $T(\Delta\setminus\{y\})$ is subgroup separable because $\Delta\setminus\{y\}$ has less vertices than $|V\Gamma|$. Ultimately, $G$ is subgroup separable as a semi-direct product of $\mathbb{Z}$ and a subgroup separable group (this follows from Theorem 2.4). ∎

### v2

# On Subgroup Separability and Membership Problems in Twisted Right-Angled Artin Groups

@@ Abstract
We characterize twisted right-angled Artin groups (T-RAAGs) that are subgroup separable using only their defining mixed graphs: such a group is subgroup separable if and only if the underlying simplicial graph contains neither induced paths nor squares on four vertices. This generalizes the results of Metaftsis–Raptis on classical right-angled Artin groups. Additionally, we show that the subgroup membership problem is decidable when the group is coherent, which occurs precisely when the defining mixed graph is chordal. We also address the rational and submonoid membership problems by exhibiting a cone-family of graphs for which the corresponding T-RAAGs have decidable rational and submonoid membership problems.
@@ 1 Introduction
In geometric group theory, the class of right-angled Artin groups (RAAGs) plays a major role, see (Charney, 2007) for a survey. RAAGs are also known as graph groups.
@@
This article deals with the same problem, with focus in another class of groups that generalises the class of RAAGs. This is the class of twisted right-angled Artin groups (T-RAAGs shortly). T-RAAGs appear in (Pride, 1986), and also in (Clancy and Ellis, 2010) – where the notion of a twisted Artin group was introduced. We refer to T-RAAGs also with the name twisted graph groups. More details regarding their normal forms and other properties can be found in the PhD thesis (Foniqi, 2022, Chapter 3), and also in (Foniqi, 2024).
@@
In Section 4 we provide our main contribution, which is the following result:
Theorem (Theorem 4.1).
Let $\Gamma$ be a mixed graph and $T(\Gamma)$ the T-RAAG based on $\Gamma$. The group $T(\Gamma)$ is subgroup separable if and only if $\Gamma$ does not contain, as induced subgraphs, mixed paths and mixed squares in $4$ vertices.
Here, in contrast with RAAGs, some of the edges in our graphs can be directed. In the case when the graph has no directed edges, we recover the result of Metaftsis and Raptis (2008) for RAAGs.
In Section 5 we discuss membership problems in T-RAAGs. Regarding membership in finitely generated subgroups we have the following:
Proposition (subsection 5.1).
Let $\Gamma$ be a mixed chordal graph. Then, $T(\Gamma)$ has decidable subgroup membership problem.
Non-separable T-RAAGs have undecidable submonoid and rational subset membership problems. On the other hand, we construct a cone-family $\mathcal{R}$ of mixed graphs (see subsection 5.2 and subsection 5.2) such that the corresponding T-RAAGs are subgroup separable, satisfying the following:
Theorem (Theorem 5.2).
Let $\Gamma$ be a mixed graph belonging to $\mathcal{R}$. Then $T(\Gamma)$ has decidable rational subset membership problem. Consequently, $T(\Gamma)$ has decidable submonoid membership problem as well.
@@ 2.2 Twisted right-angled Artin groups
To define T-RAAGs we use mixed graphs, which are similar to simplicial graphs but directed edges are allowed. We define mixed graphs using simplicial graphs as an underlying structure.
@@ 2.4 Reidemeister-Schreier procedure
Then, $Y=\{tx(\overline{tx})^{-1}\mid t\in T,\;x\in X,\;tx\not\in T\}$ gives a set of generators for $H$.
@@ 3 Subgroup separability for RAAGs
Summarising the results above, we obtain the other direction of the goal as well, as desired.
@@
Word problem is one of the fundamental algorithmic questions in algebra; for a group $G$, generated by a set $A$, the word problem asks whether $w\in(A\cup A^{-1})^{\ast}$ is equal to $1_{G}$ in $G$. This problem is solved for T-RAAGs in (Foniqi, 2022, 2024).
A more general problem is the subgroup membership problem (know also as the generalised word problem) for a group $G$, which asks for the existence of an algorithm that takes as input an element $g\in G$ and a finitely generated subgroup $H\leqslant G$ and outputs a YES – NO answer whether $g$ belongs to $H$ or not. This is still open for RAAGs, hence for T-RAAGs as well. There is a useful connection between being subgroup separable and having decidable subgroup membership problem.
@@
In (Mikhailova, 1966) it is shown that $A(C_{4})=F_{2}\times F_{2}$ contains a subgroup where membership is undecidable.
@@ 4 Going from RAAGs to T-RAAGs
We will answer one part of this question by applying the normal forms in T-RAAGs coming from (Foniqi, 2022, 2024), which yield the following two results:
@@
[…]
The result of this corollary follows also from (Pride, 1986, Theorem 2), because T-RAAGs are ‘generalized’ Artin groups with their convention, and they satisfy the $C(4)$ condition’. Moreover, paths on $4$ vertices, or a square on $4$ vertices, are triangle-free, so they satisfy their $T(4)$ condition’ as well. Therefore if ${\Gamma}$ contains a graph $\Delta$ which is a path on $4$ vertices, or a square on $4$ vertices, then $T(\Gamma)$ contains $T(\Delta)$ by section 4; by (Pride, 1986, Theorem 2) the squares of vertices of $\Delta$ generate a RAAG isomorphic to $A(\overline{\Delta})$, in which case $T(\Gamma)$ is not subgroup separable.
@@ 4.1 Subgroup Separability for T-RAAGs
By induction, the group $T(\Delta\setminus\{y\})$ is subgroup separable because $\Delta\setminus\{y\}$ has less vertices than $|V\Gamma|$. Ultimately, $G$ is subgroup separable as a semi-direct product of $\mathbb{Z}$ and a subgroup separable group (this follows from subsection 2.3). ∎
5 Membership Problem in T-RAAGs
A decision problem is a YES – NO question on a countable infinite set of inputs. A decision problem is called decidable if there is an algorithm which takes an input, terminates after a finite amount of time, and correctly answers the question by YES or NO. If such an algorithm does not exist, we refer to the decision problem as undecidable.
The word problem and the subgroup membership problem, introduced earlier, are both instances of decision problems. Natural generalizations of these classical problems include the submonoid membership problem and the rational subset membership problem, which we will define and discuss in more detail in Subsection 5.2.
5.1 Subgroup Membership Problem in T-RAAGs
The subgroup membership problem remains open for RAAGs, and hence it is open for T-RAAGs as well. However, subgroup separable T-RAAGs have a decidable subgroup membership problem (Mal’cev, 1958).
On the other hand, subgroup non-separability does not imply undecidability of the subgroup membership problem: for example, $A(P_{4})$ is not subgroup separable, yet has a decidable subgroup membership problem.
[…]
Let $\Gamma$ be a simplicial chordal graph. Then, $A(\Gamma)$ has decidable subgroup membership problem.
In contrast, the RAAG $A(C_{4})=F_{2}\times F_{2}$ contains a subgroup with undecidable membership (Mikhailova, 1966); the graph $C_{4}$ is not chordal.
RAAGs defined by chordal graphs are precisely the coherent ones (Droms, 1987).
[…]
Coherent RAAGs have decidable subgroup membership problem.
For a finite mixed graph $\Gamma$ we will say that $\Gamma$ is chordal if $\overline{\Gamma}$ is chordal.
One also has that T-RAAGs based on chordal graphs are exactly the coherent ones (see [Blumer, Foniqi, Quadrelli]).
The main goal of this secion, which uses (Kapovich, Weidmann, and Myasnikov, 2005, Theorem 1.1), is to show
[…]
Let $\Gamma$ be a mixed chordal graph. Then, $T(\Gamma)$ has decidable subgroup membership problem.
Proof.
Let $\Gamma$ be chordal. In the case when $\Gamma$ is complete, the group $T(\Gamma)$ is virtually free abelian; indeed, the subgroup generated by squares of the vertices of $\Gamma$ is free abelian, and of finite index. Threfore, $T(\Gamma)$ has decidable subgroup membership problem.
If $\Gamma$ is not complete, then there are two proper subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ with $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and $\Gamma_{1}\cap\Gamma_{2}=C$, with $C$ complete. Here one has a splitting:
$T(\Gamma)=T(\Gamma_{1})*_{T(C)}T(\Gamma_{2}).$
As $C$ is complete, $T(C)$ is virtually free abelian.
Now, by the induction hypothesis, both $T(\Gamma_{1})$, and $T(\Gamma_{2})$ are iterated amalgamated products or graphs of groups - with virtually abelian vertex groups (and edge groups). Now the result follows from induction and (Kapovich, Weidmann, and Myasnikov, 2005, Theorem 1.1). ∎
5.2 Submonoids and Rational Subsets
Let $M$ be a monoid generated by a finite set $A$, and let $\phi:A^{*}\to M$ be the canonical homomorphism.
The submonoid membership problem asks: given a finite set $S\subseteq A^{*}$ and a word $w\in A^{*}$, does $\phi(w)\in\phi(S^{*})$? This problem is well-defined for groups as well, using monoid generators $A\cup A^{-1}$.
A subset $L\subseteq M$ is rational if $L=\phi(R)$ for some regular language $R\subseteq A^{*}$. Rational subsets generalize finitely generated submonoids.
The rational subset membership problem asks: given a rational subset $R\subseteq M$ and a word $w\in A^{*}$, is $\phi(w)\in R$? This applies to groups viewed as monoids.
[…]
Decidability of rational subset membership implies decidability of submonoid membership.
These problems are preserved under passing to submonoids. Notably, a group may have decidable submonoid membership but undecidable rational subset membership (Bodart, 2024).
Theorem 5.1 (Lohrey and Steinberg, 2008).
A RAAG $A(\Gamma)$ has decidable submonoid and rational subset membership problems if and only if $\Gamma$ contains neither $P_{4}$ nor $C_{4}$ as induced subgraphs.
It is clear that if $\Gamma$ is a mixed graph of non-elementary type, then $T(\Gamma)$ has undecidable submonoid membership problem, as it contains $A(P_{4})$ or $A(C_{4})$.
On the other hand, the situation is not clear for elementary mixed graphs. But, there is an interesting sub-family of elementary T-RAAGs for which the rational subset membership problem is decidable; the construction of this family will be given below and it is motivated from the proof of Theorem 4.1.
[…]
Let $\Gamma=(V,E)$ be a mixed graph. The $\Gamma$-cone $\nabla(\Gamma)$, with tip $w$, is a mixed graph with vertex set $V\sqcup\{w\}$, such that $\Gamma$ is an induced subgraph, and every vertex $v$ of $\Gamma$ is connected to $w$ with edges of the type $[v,w]$ or $[v,w\rangle$.
[…]
Let $\mathcal{R}$ be the class of mixed graphs of elementary type, constructed by starting with a single vertex, and iterating through disjoint unions and $\Gamma$-cones.
Theorem 5.2.
Let $\Gamma$ be a mixed graph belonging in $\mathcal{R}$. $T(\Gamma)$ has decidable rational subset membership problem. Consequently, $T(\Gamma)$ has decidable submonoid membership problem.
Proof.
We proceed by induction on $n=|V\Gamma|$.
If $n=1$, then $T(\Gamma)\simeq\mathbb{Z}$, and the result is immediate.
For $n=2$, there are three possibilities:
  • (i)
    $T(\Gamma)\simeq\mathbb{Z}\ast\mathbb{Z}$, which has a decidable rational subset membership problem, as a free product of groups that do.
  • (ii)
    $T(\Gamma)\simeq\mathbb{Z}\times\mathbb{Z}$, which has a decidable rational subset membership problem, being a direct product of a group that does with $\mathbb{Z}$.
  • (iii)
    $T(\Gamma)\simeq\mathbb{Z}\rtimes\mathbb{Z}$, which also has a decidable rational subset membership problem, as a finite extension of $\mathbb{Z}^{2}$.
Now assume that $n>2$ and that the result holds for graphs in $\mathcal{R}$ with less vertices.
If $\Gamma$ is disconnected, then $\Gamma=\Gamma_{1}\sqcup\Gamma_{2}$, and so $T(\Gamma)=T(\Gamma_{1})\ast T(\Gamma_{2})$ and the result holds by induction and the closere of decidability of rational subset membership prblem under free products.
So, assume that $\Gamma$ is connected. By the construction of $\mathcal{R}$, we know that there is a vertex $w$ connected with every other vertex $v\in V\Gamma\setminus\{w\}$, by edges of the type $[v,w]$ or $[v,w\rangle$. Since graphs in $\mathcal{R}$ are mixed elementary as well, we can apply the proof of Theorem 4.1. In particular, our graph looks like the one in Figure 1, without outgoing edges, hence it looks graphically as below:
$w$$b_{1}$$b_{2}$$b_{n}$$c_{l}$$c_{1}$$c_{l}$
Now, as in the proof of Theorem 4.1, the subgroup:
$G=\langle w^{2},\operatorname{Link}(w)\rangle=\langle w^{2},b_{1},\ldots,b_{n},c_{1},\ldots,c_{l}\rangle.$
is a T-RAAG over the mixed graph
$w^{2}$$b_{1}$$b_{2}$$b_{n}$$c_{l}$$c_{1}$$c_{l}$
In particular, $G=T(\Gamma\setminus\{w\})\times\mathbb{Z}$, which has decidable rational subset membership problem, by iunduction and the closere under direct product with $\mathbb{Z}$.
On the other hand, $G$ is a subgroup of index $2$ in $T(\Gamma)$, and so $T(\Gamma)$ has decidable rational subset membership problem, by closere under finite extensions. ∎